# FARM: Functional Group-Aware Representations for Small Molecules

### v2

@@ 3.4 Molecular Structure Integration via Contrastive Learning
To integrate FG-enhanced SMILES representations with molecular structure information, we employ contrastive learning to align these representations with FG graph embeddings. This method captures both the sequential and topological aspects of molecular structures, allowing the model to develop a unified representation that integrates chemical context with overall molecular architecture.
@@
Specifically, given a positive pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})$, where $\mathbf{h}_{\text{MLM}}$ is the atom-level representation derived from a pretrained BERT model and $\mathbf{h}_{\text{pos}}$ is the corresponding structure representation from a graph neural network (GNN), and a negative pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})$, where $\mathbf{h}_{\text{neg}}$ is a augmented FG-graph, the contrastive loss can be written as:
$\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cosine\_similarity}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cosine\_similarity}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$
where $\gamma$ is the margin parameter and $N$ is the number of training examples (or contrastive pairs). The final objective function for integration is:
$\mathcal{L}_{\text{Integration}}=\lambda_{\text{MLM}}\cdot\mathcal{L}_{\text{MLM}}+\lambda_{\text{CL}}\cdot\mathcal{L}_{\text{CL}}$
where $\lambda_{\text{MLM}}$ and $\lambda_{\text{CL}}$ are hyperparameters that control the contribution of each loss to the overall objective. Details on the implementation of contrastive training can be found in the Appendix D.4
@@ 4 Pre-training Data Collection and Diversity Assessment
The performance and generalizability of a machine learning model is heavily dependent on the quality and diversity of its training data. Given the vastness of chemical space, it is vital that the training data represents a sufficiently representative subset of this space. Traditional approaches assess the diversity of datasets based on criteria such as the number of chemical elements and the number of atoms per molecule. However, to the best of our knowledge, no previous studies have systematically evaluated the diversity of large chemical datasets like ZINC (Irwin & Shoichet, 2005) or ChEMBL (Gaulton et al., 2011) in terms of functional groups. This gap is significant, as functional groups provide deep insights into the structural and functional complexity of molecules.
In this work, we introduce a novel criterion to assess dataset diversity by using the size of the FG-enhanced SMILES lexicon as a metric. By analyzing the size of this lexicon, we can assess whether the dataset captures a comprehensive range of chemical functionalities, which is crucial for building robust molecular foundation models.
Table 1 presents the size of the FG-enhanced SMILES lexicon for the ZINC15 and ChEMBL25 datasets, as well as the dataset we collected for training our foundation model. The table shows that despite its widespread use for training foundation models in small molecule representation, ZINC15 is significantly less diverse compared to ChEMBL25. This is largely due to the fact that ZINC15 is primarily designed to include molecules that adhere to Lipinski’s Rule of Five for drug-likeness, excluding more exotic or less common elements that are less relevant to pharmaceutical chemistry. This limited diversity may negatively impact the model’s performance on out-of-distribution datasets, which contain a broader range of atom types and functional groups not present in ZINC15. In contrast, ChEMBL25 is far more diverse and thus better suited for training foundation models that can be fine-tuned for various downstream tasks. Based on this insight, we collected a dataset that includes the entire ChEMBL25 database, libraries from chemical drug suppliers, and a subset of ZINC15 to ensure a more comprehensive coverage of chemical space. Detailed information about the collected data can be found in the Appendix A.
@@
Database #molecules #atom types Length $Min\_frequency=1$ $Min\_frequency=5$
@@
5 Experimental Results
Data and splits. We consider 12 benchmark tasks in the MoleculeNet dataset (Wu et al., 2018). Following previous works, the data is split using a scaffold split into training, validation, and test sets with an 8:1:1 ratio, ensuring fair and consistent evaluation across all models.
Baselines. We consider works that incorporate functional group information to enhance representation learning (Zhang et al., 2020; 2021; Yang et al., 2022; Wang et al., 2023; Han et al., 2023; Li et al., 2023; Zang et al., 2023; Wang et al., 2022b; Xia et al., 2023; Nguyen et al., 2024b), as well as other approaches that utilize masked atom prediction as a self-supervised training task (Rong et al., 2020; Hu et al., 2019; Liu et al., 2021; Fang et al., 2022; Zhou et al., 2023).
Table 2 and 3 present the performance of FARM alongside other baseline models on seven MoleculeNet classification and five regression tasks, respectively. FARM consistently outperforms baseline models on various benchmarks, demonstrating its robustness and versatility. Its strong performance across various classification and regression tasks indicates that FARM is a highly effective pretrained model, well-suited for a broad range of downstream tasks in molecular property prediction.
Table 2: Performance comparison of FARM and baseline models on MoleculeNet classification tasks. The first 10 models incorporate FG (functional group) information to enhance representation learning. Performance is evaluated using by ROC-AUC.
@@ 5 Experimental Results
Dataset BBBP Tox21 SIDER ClinTox BACE MUV HIV
#tasks 1 12 27 2 1 17 1
#samples 2,039 7,831 1,427 1,478 1,513 93,807 41,127
@@
MICRO (Zhang et al., 2020) 84.4 $\pm$ 1.1 77.0 $\pm$ 0.8 56.7 $\pm$ 0.9 77.0 $\pm$ 2.0 77.2 $\pm$ 2.0 - 75.1 $\pm$ 1.1
MGSSL (Zhang et al., 2021) 69.7 $\pm$ 0.9 76.5 $\pm$ 0.3 61.8 $\pm$ 0.8 80.7 $\pm$ 2.1 79.1 $\pm$ 0.9 78.7 $\pm$ 1.5 78.8 $\pm$ 1.2
MoleOOD (Yang et al., 2022) 71.0 $\pm$ 0.8 - 63.4 $\pm$ 0.7 - 84.3 $\pm$ 1.1 - 79.4 $\pm$ 0.5
MCM (Wang et al., 2023) 90.0 $\pm$ 3.1 80.2 $\pm$ 1.5 62.7 $\pm$ 2.8 65.5 $\pm$ 1.4 82.0 $\pm$ 5.5 - -
HimGNN (Han et al., 2023) 92.8 $\pm$ 2.7 80.7 $\pm$ 1.7 64.2 $\pm$ 2.3 91.7 $\pm$ 3.0 85.6 $\pm$ 3.4 - -
FG-BERT (Li et al., 2023) 70.2 $\pm$ 0.9 78.4 $\pm$ 0.8 64.0 $\pm$ 0.7 83.2 $\pm$ 1.6 84.5 $\pm$ 1.5 75.3 $\pm$ 2.4 77.4 $\pm$ 1.0
HiMol (Zang et al., 2023) 71.3 $\pm$ 0.6 76.0 $\pm$ 0.2 62.5 $\pm$ 0.3 70.6 $\pm$ 2.1 84.6 $\pm$ 0.2 - -
Mole-BERT (Xia et al., 2023) 71.9 $\pm$ 1.6 76.8 $\pm$ 0.5 62.8 $\pm$ 1.1 78.9 $\pm$ 3.0 80.8 $\pm$ 1.4 78.6 $\pm$ 1.8 78.2 $\pm$ 0.8
MolCLR (Wang et al., 2022b) 73.3 $\pm$ 1.0 74.1 $\pm$ 5.3 61.2 $\pm$ 3.6 89.8 $\pm$ 2.7 82.8 $\pm$ 0.7 78.9 $\pm$ 2.3 77.4 $\pm$ 0.6
GLAD (Nguyen et al., 2024b) 80.4 $\pm$ 1.5 - 64.7 $\pm$ 1.8 87.3 $\pm$ 1.2 85.7 $\pm$ 0.9 - -
N-GRAM (Hu et al., 2019) 70.8 $\pm$ 1.5 78.7 $\pm$ 0.4 62.7 $\pm$ 0.8 72.6 $\pm$ 1.5 84.5 $\pm$ 0.7 81.3 $\pm$ 2.1 79.9 $\pm$ 0.7
GROVER (Rong et al., 2020) 86.8 $\pm$ 2.2 80.3 $\pm$ 2.0 61.2 $\pm$ 2.5 70.3 $\pm$ 13.7 82.4 $\pm$ 3.6 67.3 $\pm$ 1.8 68.2 $\pm$ 1.1
GraphMVP (Liu et al., 2021) 72.4 $\pm$ 1.6 75.9 $\pm$ 0.5 63.1 $\pm$ 0.4 79.1 $\pm$ 2.8 81.2 $\pm$ 0.9 77.7 $\pm$ 0.6 77.0 $\pm$ 1.2
GEM (Fang et al., 2022) 88.8 $\pm$ 0.4 78.1 $\pm$ 0.4 63.2 $\pm$ 1.5 90.3 $\pm$ 0.7 87.9 $\pm$ 1.1 75.3 $\pm$ 1.5 81.3 $\pm$ 0.3
UniMol (Zhou et al., 2023) 72.9 $\pm$ 0.6 79.6 $\pm$ 0.5 65.9 $\pm$ 1.3 91.9 $\pm$ 1.8 85.7 $\pm$ 0.2 82.1 $\pm$ 1.3 82.8 $\pm$ 0.3
FARM (Ours) 93.3 $\pm$ 0.2 80.8 $\pm$ 1.1 65.9 $\pm$ 0.7 82.2 $\pm$ 0.7 89.6 $\pm$ 0.4 82.7 $\pm$ 2.1 83.5 $\pm$ 0.5
@@
6 Conclusion, Limitations and Future Work
In summary, FARM demonstrates robust performance across various MoleculeNet classification tasks, outperforming or matching baseline models. The integration of functional group information and the alignment of FG-enhanced SMILES representations with FG graph embeddings through contrastive learning significantly enhance its effectiveness. This approach underscores FARM’s versatility and strength as a pre-trained model, capable of improving molecular structure understanding and predictive accuracy for a wide range of downstream tasks.
While FARM shows strong performance, there are two main limitations that should be addressed in future work. First, the current model does not incorporate a full 3D molecular representation, which is critical for capturing stereochemistry and spatial configurations that affect molecular properties. Incorporating 3D information Yan et al. (2024) could further enhance the model’s predictions. Second, the model faces challenges when dealing with rare fused ring systems due to out-of-vocabulary issues. A potential solution to this limitation is to extend the training dataset, covering a broader portion of chemical space to include more diverse and complex molecular structures.
@@ A.1 Training data
A.2 Downstream Tasks Data
@@ Appendix C FG Knowledge Graph
Table 6: Key relations defined in the FG knowledge graph.
@@ D.1 Training Masked Language Model for SMILES Representation
We trained the BERT model using Hugging Face (Wolf et al., 2020) on the masked molecule prediction task with both conventional SMILES and FG-enhanced SMILES from our collected dataset. To assess the impact of different masking percentages, we trained BERT models with masking percentages of 0.15, 0.25, 0.35, 0.45, and 0.55. The models were then evaluated on seven MoleculeNet tasks, including three classification tasks and four regression tasks, to determine the optimal masking percentage. The results, presented in Table 7, indicate that a masking percentage of 0.35 yields the best performance across the considered downstream tasks.
@@
Additional details of the training setup include training the BERT model on 20 million SMILES for 15 epochs using two NVIDIA Tesla V100 GPUs. The learning rate was set to $51e-$ 5 , with a batch size of 128, and model checkpoints were saved after every 10,000 batches. This setup was also applied to the baseline model, which used conventional SMILES for comparison.

### v3

@@ 3.4 Molecular Structure Integration via Contrastive Learning
To integrate FG-enhanced SMILES representations with molecular structure information, we employ contrastive learning to align these representations with FG graph embeddings. This method captures both the atom-level and topological aspects of molecular structures, allowing the model to develop a unified representation that integrates chemical context with overall molecular architecture.
@@
Specifically, given a positive pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})$, where $\mathbf{h}_{\text{MLM}}$ is the atom-level representation derived from a pretrained BERT model and $\mathbf{h}_{\text{pos}}$ is the corresponding structure representation from a graph neural network (GNN), and a negative pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})$, where $\mathbf{h}_{\text{neg}}$ is a augmented FG-graph, the contrastive loss can be written as: $\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$.
where $\gamma$ is the margin parameter and $N$ is the number of training examples (or contrastive pairs). The final objective function for integration is: $\mathcal{L}_{\text{Integration}}=\lambda_{\text{MLM}}\cdot\mathcal{L}_{\text{MLM}}+\lambda_{\text{CL}}\cdot\mathcal{L}_{\text{CL}}$, where $\lambda_{\text{MLM}}$ and $\lambda_{\text{CL}}$ are hyperparameters that control the contribution of each loss to the overall objective. Details on the implementation of contrastive training can be found in the Appendix D.4.
@@ 4 Pre-training Data Collection and Diversity Assessment
The performance and generalizability of a machine learning model is heavily dependent on the quality and diversity of its training data. Given the vastness of chemical space, it is vital that the training data represents a sufficiently representative subset of this space. Traditional approaches assess the diversity of datasets based on criteria such as the number of chemical elements and the number of atoms per molecule. However, to the best of our knowledge, no previous studies have systematically evaluated the diversity of large chemical datasets like ZINC (Irwin and Shoichet, 2005) or ChEMBL (Gaulton et al., 2011) in terms of functional groups. This gap is significant, as functional groups provide deep insights into the structural and functional complexity of molecules.
In this work, we introduce a novel criterion to assess dataset diversity by using the size of the FG-enhanced SMILES lexicon as a metric. By analyzing the size of this lexicon, we can assess whether the dataset captures a comprehensive range of chemical functionalities, which is crucial for building robust molecular foundation models. Table 1 presents the size of the FG-enhanced SMILES lexicon for the ZINC15 and ChEMBL25 datasets, as well as the dataset we collected for training our foundation model. The table shows that despite its widespread use for training foundation models in small molecule representation, ZINC15 is significantly less diverse compared to ChEMBL25. This is largely due to the fact that ZINC15 is primarily designed to include molecules that adhere to Lipinski’s Rule of Five for drug-likeness, excluding more exotic or less common elements that are less relevant to pharmaceutical chemistry. This limited diversity may negatively impact the model’s performance on out-of-distribution datasets, which contain a broader range of atom types and functional groups not present in ZINC15. In contrast, ChEMBL25 is far more diverse and thus better suited for training foundation models that can be fine-tuned for various downstream tasks. Based on this insight, we collected a dataset that includes the entire ChEMBL25 database, libraries from chemical drug suppliers, and a subset of ZINC15 to ensure a more comprehensive coverage of chemical space. Detailed information about the collected data can be found in the Appendix A.
@@
[…]
@@
Table 2: Performance comparison of FARM and baseline models on MoleculeNet classification tasks. The first 10 models incorporate FG (functional group) information to enhance representation learning.
@@
Dataset BBBP Tox21 ToxCast SIDER ClinTox BACE MUV HIV
#tasks 1 12 617 27 2 1 17 1
#samples 2,039 7,831 8,575 1,427 1,478 1,513 93,807 41,127
@@
MICRO (Zhang et al., 2020) 84.4 $\pm$ 1.1 77.0 $\pm$ 0.8 65.2 $\pm$ 0.8 56.7 $\pm$ 0.9 77.0 $\pm$ 2.0 77.2 $\pm$ 2.0 - 75.1 $\pm$ 1.1
MGSSL (Zhang et al., 2021) 69.7 $\pm$ 0.9 76.5 $\pm$ 0.3 64.1 $\pm$ 0.7 61.8 $\pm$ 0.8 80.7 $\pm$ 2.1 79.1 $\pm$ 0.9 78.7 $\pm$ 1.5 78.8 $\pm$ 1.2
MoleOOD (Yang et al., 2022) 71.0 $\pm$ 0.8 - - 63.4 $\pm$ 0.7 - 84.3 $\pm$ 1.1 - 79.4 $\pm$ 0.5
MCM (Wang et al., 2023) 90.0 $\pm$ 3.1 80.2 $\pm$ 1.5 - 62.7 $\pm$ 2.8 65.5 $\pm$ 1.4 82.0 $\pm$ 5.5 - -
HimGNN (Han et al., 2023) 92.8 $\pm$ 2.7 80.7 $\pm$ 1.7 - 64.2 $\pm$ 2.3 91.7 $\pm$ 3.0 85.6 $\pm$ 3.4 - -
FG-BERT (Li et al., 2023) 70.2 $\pm$ 0.9 78.4 $\pm$ 0.8 63.3 $\pm$ 0.8 64.0 $\pm$ 0.7 83.2 $\pm$ 1.6 84.5 $\pm$ 1.5 75.3 $\pm$ 2.4 77.4 $\pm$ 1.0
HiMol (Zang et al., 2023) 71.3 $\pm$ 0.6 76.0 $\pm$ 0.2 - 62.5 $\pm$ 0.3 70.6 $\pm$ 2.1 84.6 $\pm$ 0.2 - -
Mole-BERT (Xia et al., 2023) 71.9 $\pm$ 1.6 76.8 $\pm$ 0.5 62.8 $\pm$ 1.1 62.8 $\pm$ 1.1 78.9 $\pm$ 3.0 80.8 $\pm$ 1.4 78.6 $\pm$ 1.8 78.2 $\pm$ 0.8
MolCLR (Wang et al., 2022b) 73.3 $\pm$ 1.0 74.1 $\pm$ 5.3 - 61.2 $\pm$ 3.6 89.8 $\pm$ 2.7 82.8 $\pm$ 0.7 78.9 $\pm$ 2.3 77.4 $\pm$ 0.6
GLAD (Nguyen et al., 2024b) 80.4 $\pm$ 1.5 - - 64.7 $\pm$ 1.8 87.3 $\pm$ 1.2 85.7 $\pm$ 0.9 - -
N-GRAM (Hu et al., 2019) 70.8 $\pm$ 1.5 78.7 $\pm$ 0.4 66.5 $\pm$ 0.3 62.7 $\pm$ 0.8 72.6 $\pm$ 1.5 84.5 $\pm$ 0.7 81.3 $\pm$ 2.1 79.9 $\pm$ 0.7
GROVER (Rong et al., 2020) 86.8 $\pm$ 2.2 80.3 $\pm$ 2.0 65.3 $\pm$ 0.5 61.2 $\pm$ 2.5 70.3 $\pm$ 13.7 82.4 $\pm$ 3.6 67.3 $\pm$ 1.8 68.2 $\pm$ 1.1
GraphMVP (Liu et al., 2021) 72.4 $\pm$ 1.6 75.9 $\pm$ 0.5 63.1 $\pm$ 0.4 63.1 $\pm$ 0.4 79.1 $\pm$ 2.8 81.2 $\pm$ 0.9 77.7 $\pm$ 0.6 77.0 $\pm$ 1.2
GEM (Fang et al., 2022) 88.8 $\pm$ 0.4 78.1 $\pm$ 0.4 69.2 $\pm$ 0.4 63.2 $\pm$ 1.5 90.3 $\pm$ 0.7 87.9 $\pm$ 1.1 75.3 $\pm$ 1.5 81.3 $\pm$ 0.3
UniMol (Zhou et al., 2023) 72.9 $\pm$ 0.6 79.6 $\pm$ 0.5 69.6 $\pm$ 0.1 65.9 $\pm$ 1.3 91.9 $\pm$ 1.8 85.7 $\pm$ 0.2 82.1 $\pm$ 1.3 82.8 $\pm$ 0.3
FARM (Ours) 93.3 $\pm$ 0.2 80.8 $\pm$ 1.1 69.9 $\pm$ 0.5 65.9 $\pm$ 0.7 82.2 $\pm$ 0.7 89.6 $\pm$ 0.4 82.7 $\pm$ 2.1 83.5 $\pm$ 0.5
@@
5 Experimental Results
Data and Splits   We consider 13 benchmark tasks in the MoleculeNet dataset (Wu et al., 2018). Following previous works, the data is split using a scaffold split into training, validation, and test sets with an 8:1:1 ratio, ensuring fair and consistent evaluation across all models.
Baselines   We consider works that incorporate functional group information to enhance representation learning (Zhang et al., 2020, 2021; Yang et al., 2022; Wang et al., 2023; Han et al., 2023; Li et al., 2023; Zang et al., 2023; Wang et al., 2022b; Xia et al., 2023; Nguyen et al., 2024b), and other approaches that utilize masked atom prediction as a self-supervised training task (Rong et al., 2020; Hu et al., 2019; Liu et al., 2021; Fang et al., 2022; Zhou et al., 2023).
Results   Table 2 and 3 present the performance of FARM alongside other baseline models on seven MoleculeNet classification tasks. FARM consistently outperforms baseline models on various benchmarks, demonstrating its robustness and versatility. Its strong performance across various classification and regression tasks indicates that FARM is a highly effective pretrained model, well-suited for a broad range of downstream tasks in molecular property prediction.
Ablation Study on FARM Component Analysis   To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The results are shown in Table 10. For more details about the baselines, please refer to appendix F. The results show that each component of FARM contributes to improved performance compared to the baseline. Combining all components leads to strong overall performance of FARM across all test tasks.
Ablation Study on Fragmentation Methods   We compare our FG-based fragmentation method with BRICS fragmentation. When treating each BRICS fragment as a functional group, the enriched SMILES vocabulary size reaches 575,612 (min freq $\leq$ 5), compared to just 14,741 with our FG-based method. This is because BRICS produces larger, less reusable fragments, resulting in a bloated vocabulary with many rare tokens—making training harder and increasing the risk of out-of-vocabulary issues. As shown in Table 4, our FG-based method consistently outperforms BRICS across all six MoleculeNet tasks under the same setup.
Table 4: Performance comparison of the model using BRICS fragmentation versus our FG-based fragmentation across six MoleculeNet tasks.
BBBP BACE HIV Average ESOL FreeSolv Average QM9
#tasks 1 1 1 1 1 3
#samples 2039 1513 41127 1128 642 133885
Metric ROC-AUC ($\uparrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
BRICS 88.87 $\pm$ 1.5 89.54 $\pm$ 1.09 70.09 $\pm$ 2.24 82.83 1.340 $\pm$ 0.235 0.823 $\pm$ 0.14 1.081 0.0085 $\pm$ 0.00034
OURS 96.23 $\pm$ 0.7 96.19 $\pm$ 0.65 82.13 $\pm$ 1.10 91.43 0.734 $\pm$ 0.039 0.308 $\pm$ 0.08 0.521 0.0038 $\pm$ 0.00014
6 Conclusion, Limitations, and Future Work
In summary, FARM demonstrates robust performance across various MoleculeNet classification tasks, outperforming or matching baseline models. The integration of functional group information and the alignment of FG-enhanced SMILES representations with FG graph embeddings through contrastive learning significantly enhance its effectiveness. While FARM shows strong performance, there is a key limitation that should be addressed in future work. The current model does not incorporate a full 3D molecular representation, which is critical for capturing stereochemistry and spatial configurations that influence molecular properties. Integrating 3D structural information Yan et al. (2024) could further improve the model’s ability to make accurate predictions.
@@ A.1 Training data
A.2 Lexicon Growth for FG-enhanced SMILES Dataset
As part of our investigation into the relationship between corpus size and vocabulary size, we conducted experiments to observe how vocabulary growth evolves as the dataset increases. Our results show that while the vocabulary size increases with the corpus size, it eventually stabilizes. Specifically, when expanding the corpus from 20 million tokens to 200 million tokens, the vocabulary grew from 14,741 to 42,364 tokens, and with a 1 billion-token corpus, the vocabulary reached a plateau of 42,391 tokens. This indicates that the lexicon tends to converge around 42.3K tokens as the corpus size increases, suggesting that further growth in corpus size may not result in substantial increases in vocabulary.
Even with the smaller vocabulary of 14,741 tokens derived from the 20 million-token dataset, the model was able to achieve excellent coverage of common tokens. In fact, the 14,741 tokens from the 20M dataset accounted for 99.997% of all tokens in the 200M dataset, with 2,932,883,078 out of 2,932,984,142 tokens already represented. This demonstrates that a relatively smaller vocabulary can effectively capture the majority of the data, even with larger corpora. Moreover, the ratio of the UNKNOWN token in downstream datasets remains consistently below 0.1
A.3 Downstream Tasks Data
@@ Appendix C FG Knowledge Graph
Table 7: Key relations defined in the FG knowledge graph. (Note: Continuous values, such as LogP and water solubility, are discretized by rounding to the nearest integer.)
@@ D.1 Training Masked Language Model for SMILES Representation
We trained the BERT model using Hugging Face [Wolf et al., 2020] on the masked molecule prediction task with both conventional SMILES and FG-enhanced SMILES from our collected dataset. To assess the impact of different masking percentages, we trained BERT models with masking percentages of 0.15, 0.25, 0.35, 0.45, and 0.55. The models were then evaluated on seven MoleculeNet tasks, including three classification tasks and four regression tasks, to determine the optimal masking percentage. The results, presented in Table 8, indicate that a masking percentage of 0.35 yields the best performance across the considered downstream tasks.
@@
Additional details of the training setup include training the BERT model on 20 million SMILES for 15 epochs using two NVIDIA Tesla V100 GPUs. The learning rate was set to $1e-5$ , with a batch size of 128, and model checkpoints were saved after every 10,000 batches. This setup was also applied to the baseline model, which used conventional SMILES for comparison.